# ThermoNeRF: Multimodal Neural Radiance Fields for Thermal Novel View Synthesis

### v1

@@ 6 Experiments
We conduct a comprehensive evaluation of the quality of rendered novel views using the ThermoScenes dataset introduced in Section 5. This section details the evaluation metrics, compares baselines, and discusses experimental results and ablation studies.
@@ 6.1 Evaluation Metrics
Temperature Metrics. We assess the accuracy of the rendered temperatures by computing the Mean Absolute Error (MAE) against the ground-truth thermal view. Given that thermal images often display uniform temperature distributions outside the regions of interest, calculating the MAE across the entire image may disproportionately emphasize the ambient background temperature. From an application standpoint, the temperature of the region of interest (ROI) is often crucial for temperature assessment [7]. Therefore, we also report the , the MAE computed over the region of interest. We use Otsu’s method [42] to determine the optimal threshold that distinguishes the region-of-interest pixels from the background in thermal images.
Image Quality Metrics. We report the image quality metrics Peak-Signal-to-Noise-Ratio (PSNR) [23] and Structural Similarity Metric (SSIM) [54] for both thermal and RGB modalities. For the RGB views, we include the Learned Perceptual Image Patch Similarity (LPIPS) [60], which is specifically designed to evaluate the human-perceived similarity for RGB images.
@@ 6.2 Baselines
We conduct experiments to evaluate the two design choices of our method, i.e. (1) the necessity of multimodality, as thermal images alone cannot accurately estimate the scene density due to lack of texture details, and (2) the decoupling of the RGB and thermal modalities, stemming from the physical independence between temperature and color information. Given that ThermoNeRF is based on Nerfacto, we define two baseline methods derived from Nerfacto: Nerfacto and Nerfacto.
Nerfacto is trained exclusively with thermal inputs (without RGB). It processes thermal images as single-channel grayscale images through Nerfacto’s standard pipeline. This baseline is designed to highlight the importance of incorporating the RGB modality.
Nerfacto takes both RGB and thermal modalities as inputs by concatenating them into four-channel images and optimizes the concatenated RGB-thermal images without employing separate MLPs for optimizing each modality. This baseline is designed to show the difference between simply concatenating the modalities versus optimizing them separately with a dedicated MLP for each.
@@ 6.4 Thermal View Synthesis
Tab. 2 provides a comparative analysis of our method’s performance for novel view synthesis in the thermal modality, across the ten scenes of the ThermoScenes dataset.
@@
Temperature Metrics. Tab. 2 reports the average per-pixel MAE and MAE of thermal views synthesis on the test set. With an average MAE of C, ThermoNeRF significantly outperforms (C) and (C). When trained on thermal images alone, fails to learn an accurate representation of the scene (as visible in Fig. 4) and has the lowest image quality and temperature estimation results. On the other hand, thanks to the concatenation of RGB and thermal information, obtains better results than . This is especially visible on the ROI with an average MAE of C, against C for . However, the estimated temperatures are impacted by the joint optimization of both modalities in the same MLP, as evidenced in Fig. 5 where the error in temperature prediction is high on the region of interest. On the contrary, across the ten scenes, the average MAE and MAE for ThermoNerf are C and C, respectively, marking a significant improvement over the second-best result of C and C.
Image Quality Metrics. ThermoNeRF achieves either similar or significantly better results than for both the PSNR and SSIM metrics, with a PSNR and SSIM of and , against and for . Using only thermal images results in the lowest PSNR and SSIM ( and respectively), highlighting the importance of using both modalities.
In conclusion, the proposed ThermoNeRF architecture brings an improvement of over 50% in temperature estimation, and improved or similar PSNR and SSIM, compared to the second-best model, across various outdoor and indoor scenes.
Metric | Method |
Trees | Avg | ||||||||||||||||||
PSNR | 23.68 | 29.25 | 23.34 | 18.50 | 20.30 | 22.80 | 10.49 | 18.08 | 23.88 | 20.91 | 21.12 | |||||||||||||||||||
29.76 | 31.80 | 22.9 | 32.70 | 20.60 | 28.47 | 29.82 | 24.3 | 27.74 | 31.48 | 27.96 | ||||||||||||||||||||
ThermoNeRF | 32.05 | 34.04 | 30.67 | 32.24 | 26.63 | 28.75 | 30.75 | 31.80 | 33.79 | 31.07 | 31.18 | |||||||||||||||||||
SSIM | 0.71 | 0.89 | 0.95 | 0.93 | 0.91 | 0.87 | 0.45 | 0.71 | 0.94 | 0.92 | 0.83 | |||||||||||||||||||
0.83 | 0.91 | 0.96 | 0.98 | 0.89 | 0.89 | 0.89 | 0.82 | 0.95 | 0.94 | 0.91 | ||||||||||||||||||||
ThermoNeRF | 0.92 | 0.94 | 0.97 | 0.98 | 0.92 | 0.88 | 0.95 | 0.96 | 0.97 | 0.94 | 0.94 | |||||||||||||||||||
MAE | 13.57 | 5.18 | 6.75 | 12.27 | 6.36 | 1.80 | 2.85 | 4.82 | 1.29 | 1.60 | 5.65 | |||||||||||||||||||
5.35 | 3.25 | 10.33 | 1.26 | 6.54 | 0.86 | 1.06 | 1.62 | 1.00 | 0.31 | 3.16 | ||||||||||||||||||||
ThermoNeRF | 2.10 | 2.76 | 3.26 | 1.57 | 1.88 | 0.66 | 0.91 | 1.28 | 0.31 | 0.25 | 1.50 | |||||||||||||||||||
MAE | 1.82 | 4.19 | 1.67 | 1.97 | 6.74 | 1.87 | 5.33 | 2.25 | 1.56 | 1.49 | 2.89 | |||||||||||||||||||
0.87 | 1.54 | 2.01 | 0.46 | 6.59 | 0.89 | 0.55 | 1.15 | 0.97 | 0.34 | 1.54 | ||||||||||||||||||||
ThermoNeRF | 0.53 | 0.71 | 0.57 | 0.29 | 2.40 | 0.76 | 0.34 | 0.27 | 0.35 | 0.33 | 0.66 |
@@
We visually compare examples of test views rendered by each of the baseline methods as well as by ThermoNeRF (see Fig. 4). We observe noisy renderings with , while Nerfacto generates sensible reconstruction of the geometry—some scenes such as the Double Robot scene being noisier than others. However, temperature predictions are influenced by the RGB modality, leading to biased temperature values and sharper edges compared to the ground truth for the Building (Spring). Moreover, we can see in the rendered thermal images, elements from the RGB view that are invisible in the ground truth thermal images—this is especially visible in the Raspberry Pi scene with the mouse and keyboard (Fig. 3(c)). This indicates that the RGB channels have influenced the temperature estimates, as both modalities are optimized within the same MLP. Similar renderings for all ten datasets are shown in the Results section of the supplementary material.
@@
In Fig. 5, where we visualize the per-pixel absolute errors in temperature prediction, we observe highest errors for all models at the edges. This is due to the ghosting effect available in thermal images where the edges are blurry. We also can observe that for Building (Spring) (top-left), the sky has more errors than the Exhibition Building scene (top-right). That is due to having a mix of clear sky and cloudy images on the Building (Spring) scene, as opposed to just a cloudy sky in the Exhibition Building scene. The inconsistencies of thermal radiation measured from the sky leading to less accurate thermal predictions. Overall, ThermoNeRF exhibits the best thermal prediction results while Nerfacto shows the worst thermal predictions as its reconstructions is very distorted. Nerfacto shows higher errors across the scenes than ThermoNeRF due to the contamination of the thermal information by the RGB as both modalities are optimised in a single MLP.
@@ 6.5 RGB View Synthesis
Trees | Avg | ||||||||||||||||||
PSNR | Nerfacto (RGB) | 18.11 | 23.00 | 24.71 | 17.97 | 19.88 | 20.02 | 19.05 | 19.36 | 22.18 | 20.37 | 20.46 | ||||||||||||||||||
ThermoNeRF | 18.16 | 21.05 | 25.69 | 16.20 | 19.30 | 19.74 | 19.02 | 19.07 | 22.81 | 18.67 | 19.97 | |||||||||||||||||||
Difference | +0.05 | -1.95 | +0.98 | -1.77 | -0.58 | -0.28 | -0.03 | -0.29 | +0.83 | -1.70 | -0.49 | |||||||||||||||||||
SSIM | Nerfacto (RGB) | 0.53 | 0.75 | 0.73 | 0.52 | 0.64 | 0.62 | 0.65 | 0.73 | 0.64 | 0.67 | 0.65 | ||||||||||||||||||
ThermoNeRF | 0.56 | 0.60 | 0.77 | 0.43 | 0.63 | 0.61 | 0.65 | 0.72 | 0.66 | 0.59 | 0.62 | |||||||||||||||||||
Difference | +0.03 | -0.15 | +0.04 | -0.09 | -0.02 | -0.01 | 0.00 | -0.01 | +0.02 | -0.08 | -0.03 | |||||||||||||||||||
LPIPS | Nerfacto (RGB) | 0.25 | 0.23 | 0.42 | 0.17 | 0.24 | 0.34 | 0.28 | 0.18 | 0.24 | 0.32 | 0.27 | ||||||||||||||||||
ThermoNeRF | 0.25 | 0.24 | 0.37 | 0.18 | 0.27 | 0.34 | 0.30 | 0.19 | 0.24 | 0.35 | 0.27 | |||||||||||||||||||
Difference | 0.00 | +0.01 | -0.08 | +0.01 | +0.03 | 0.00 | +0.02 | +0.01 | 0.00 | +0.03 | 0.00 |
To assess novel view synthesis with ThermoNeRF in the RGB modality, we investigate whether there is a significant degradation in the quality of RGB views compared with the standard Nerfacto trained only on the RGB inputs. Tab. 2 summarises the results, showing little to no degradation in the quality of the RGB views. Indeed, the differences between the PSNR, SSIM and LPIPS of Nerfacto and ThermoNeRF in RGB are , and , respectively.
6.6 Ablation Study
In this section, we conduct an ablation study to evaluate the training strategy of our proposed method. Since ThermoNeRF decouples the MLPs for RGB and thermal modalities, the ablation experiments are designed to assess our fully joint training strategy—especially, as opposed to sequential training strategies in which the model is first trained on RGB data and then fine-tuned using thermal data. We compare three different strategies, trained over ThermoScenes:
Strategy A initially trains the model using RGB data for 30k iterations, and then fine-tunes the model using thermal data for 30k more iterations. Hence, is frozen for the first half of the training period, directing the density optimisation to rely entirely on the RGB modality. In the second half of the training, is frozen, and the remaining MLPs are trained for 30k iterations.
Strategy B never uses thermal information to train , which relies solely on RGB data. Similar to the structure of strategy A, strategy B involves freezing throughout the first half of the training period. However, in the second half, both and are frozen.
Strategy C initially trains using only RGB data, before training RGB and thermal modalities together. For strategy C, remains frozen during the first half of the training, similar to the previous cases. However, unlike the earlier strategies, this case involves joint fine-tuning of all modalities during the latter 30k iterations.
We compare the three training strategies with our proposed ThermoNeRF strategy, which jointly trains all modalities for 30k iterations.
Tab. 4 summarises the strategies explored and presents the results for the thermal and RGB modalities with each training strategy, averaged across the ten scenes of ThermoScenes. We report average PSNR, SSIM, and MAE for thermal views on the test sets, as well as PSNR, SSIM, and LPIPS for RGB views using the same test poses. Refer to Ablations section of the appendix for per-scene results. The results demonstrate that ThermoNeRF is the most effective training strategy. Additionally, the findings indicate that density cannot be effectively learned on one modality and then transferred to another without additional joint fine-tuning. More specifically, strategy A has an of 7.38°C, 6.11°C more than ThermoNeRF. This highlights that relying solely on thermal images is insufficient, even if the density has already been optimised with the RGB modality. Optimising thermal information alone without re-optimising the density, as done in strategy B, leads to increasing the thermal prediction error in the ROI by 7.03°C compared to ThermoNeRF, on average. On the other hand, strategy’s is 2.02°C, an increase in error of 0.75°C compared to ThermoNeRF. While strategy C demonstrates that joint training outperforms other disjoint training strategies, it does not validate the effectiveness of first training using the RGB data followed by multimodal fine-tuning.
@@ 7 Conclusion
In this work, we propose ThermoNeRF, a novel multimodal approach leveraging Neural Radiance Fields (NeRF) for rendering novel RGB and thermal views of a scene. Additionally, we have curated a new dataset specifically designed for RGB+thermal scene reconstruction.
To compensate for the absence of texture in thermal images, our approach leverages paired RGB and thermal data throughout the training process. While the scene density is effectively learned from RGB and thermal inputs by a shared MLP, distinct networks are applied to estimate color and temperature information. Our experimental findings reveal that ThermoNeRF excels in synthesizing thermal images, achieving an average mean absolute error of just 1.5°C, representing more than 50% improvement over a baseline concatenating the RGB and thermal modalities. Moreover, we show our multimodal approach is not detrimental to the quality of the resulting RGB views. Finally, an ablation study demonstrates that ThermoNeRF’s fully-joint multimodal training is more effective than several alternative disjoint training strategies.
An compelling avenue for future research is exploring the adaptability of ThermoNeRF across different frameworks. Although our present work builds upon Nerfacto, investigating ThermoNeRF’s compatibility with recent anti-aliasing NeRF variants, such as ZIP-NeRF [4], presents an interesting opportunity to assess its flexibility and potential for broader application.
Furthermore, a potential limitation of our method is its dependence on paired RGB and thermal images. Collecting such paired RGB-thermal images can pose practical challenges, especially depending on the available hardware. While RGB+thermal cameras—similar to the one used in our experiments—are accessible and cost-effective, using separate RGB and thermal cameras requires careful calibration to align these modalities accurately. Future work will focus on developing a method that allows to train ThermoNeRF using unpaired RGB and thermal image data.

### v2

@@ 6 Experiments
| Metric | Method |
Laptop | Trees | Avg | |||||||||||
| PSNR | 23.68 | 29.25 | 23.34 | 18.50 | 10.49 | 18.08 | 18.28 | 20.91 | 20.32 | ||||||||||||
| 29.76 | 31.80 | 22.9 | 32.70 | 29.82 | 24.30 | 24.28 | 31.48 | 28.38 | |||||||||||||
| ThermoNeRF | 32.05 | 34.04 | 30.67 | 32.24 | 30.75 | 31.80 | 25.40 | 31.07 | 31.00 | ||||||||||||
| SSIM | 0.71 | 0.89 | 0.95 | 0.93 | 0.45 | 0.71 | 0.65 | 0.92 | 0.78 | ||||||||||||
| 0.83 | 0.91 | 0.96 | 0.98 | 0.89 | 0.82 | 0.73 | 0.94 | 0.88 | |||||||||||||
| ThermoNeRF | 0.92 | 0.94 | 0.97 | 0.98 | 0.95 | 0.96 | 0.74 | 0.94 | 0.92 | ||||||||||||
| MAE ∘C | 13.57 | 5.18 | 6.75 | 12.27 | 2.85 | 4.82 | 11.33 | 1.60 | 7.30 | ||||||||||||
| 5.35 | 3.25 | 10.33 | 1.26 | 1.06 | 1.62 | 0.54 | 0.31 | 2.96 | |||||||||||||
| ThermoNeRF | 2.10 | 2.76 | 3.26 | 1.57 | 0.91 | 1.28 | 0.43 | 0.25 | 1.57 | ||||||||||||
| MAE ∘C | 1.82 | 4.19 | 1.67 | 1.97 | 5.33 | 2.25 | 0.54 | 1.49 | 2.41 | ||||||||||||
| 0.87 | 1.54 | 2.01 | 0.46 | 0.55 | 1.15 | 0.33 | 0.34 | 0.91 | |||||||||||||
| ThermoNeRF | 0.53 | 0.71 | 0.57 | 0.29 | 0.34 | 0.27 | 0.28 | 0.33 | 0.41 |
We conduct a comprehensive evaluation of rendered novel views using ThermoScenes. This section details the evaluation metrics, compares baselines, and discusses results.
@@ 6.1 Evaluation Metrics
Temperature Metrics. We assess the accuracy of the rendered temperatures by calculating the Mean Absolute Error (MAE). Given that thermal images often display uniform temperature distributions outside the regions of interest (ROI), calculating the MAE across the entire image may disproportionately emphasize the ambient background temperature. From an application standpoint, the temperature of the ROI (e.g., the building facade) is crucial for temperature assessment [53]. Therefore, we also report the , the MAE computed over the region of interest. We use Otsu’s method [54] to determine the optimal threshold that distinguishes the ROI pixels from the background in thermal images. We also report the standard image quality metrics to evaluate the reconstruction quality for RGB and thermal images.
Image Quality Metrics. We report the commonly used image quality metrics Peak-Signal-to-Noise-Ratio (PSNR) [55] and Structural Similarity Metric (SSIM) [56] for both thermal and RGB modalities. For the RGB views, we also include the Learned Perceptual Image Patch Similarity (LPIPS) [57]; since the LPIPS is specifically designed to evaluate the human-perceived similarity for RGB images, we do not include it for thermal images evaluation.
Note that most works in RGB+thermal novel view synthesis have focused on evaluating only image quality metrics, either solely for the RGB modality [18, 40]—highlighting the visual improvement in RGB brought by the use of thermal images—or on both RGB and thermal modalities [16, 17]. In both cases, the visual quality of the generated images is being evaluated. However, the differences between the estimated and ground-truth temperature values is not measured nor analyzed. In the present work, we rather focus on temperature metrics—while ensuring good visual quality—as accurately reconstructing the surface temperature is crucial for real-world applications such as building envelope analysis for retrofit purposes.
@@ 6.2 Baselines
We conduct experiments to evaluate (1) the necessity of multimodality, as thermal images alone cannot accurately estimate the scene density due to lack of texture details, and (2) the decoupling of the RGB and thermal modalities, stemming from the physical independence between temperature and color information. Given that ThermoNeRF is based on Nerfacto, we define two baseline methods derived from Nerfacto: Nerfacto and Nerfacto.
Nerfacto is trained with thermal inputs only and processes thermal images as single-channel grayscale images through Nerfacto’s standard pipeline. It highlights the importance of incorporating the RGB modality.
Nerfacto takes both RGB and thermal modalities as inputs by concatenating them into four-channel images and optimizes the concatenated RGB-thermal images without employing separate MLPs for optimizing each modality. Tests against Nerfacto highlights the importance of disjoint optimization of the modalities.
@@ 6.3 Implementation Details
We use 2-layer MLP with width 64 and 1 output dimension for MLP and MLP. For MLP, we use 3-layer MLP with width 64 and 3 output dimensions. We use ReLU activation function for all MLPs and Sigmoid function for the head of MLP and MLP. We use the default NeRFstudio [58] settings for the hash encoder which has a minimum and maximum resolution of 16 and 1024 respectively and 2 features per level. Additionally, we use the default settings of the spherical harmonic encoding with 4 spherical harmonic levels. All details are available in our code repository666https://github.com/Schindler-EPFL-Lab/thermo-nerf.
@@ 6.4 Thermal View Synthesis
Table 3 and Table 2 provide a comparative analysis of our method’s performance for novel view synthesis for the thermal modality on the ThermoScenes dataset.
@@
With an average MAE of C on building scenes and C on other scenes, ThermoNeRF significantly outperforms (C and C) and (C and C). When trained on thermal images alone, fails to learn an accurate representation of the scene (as visible in Fig. 4(b)) and has the lowest image quality and temperature estimation results. On the other hand, thanks to the concatenation of RGB and thermal information, obtains better results than : MAE of C, against C for on building scenes, and C against C on other scenes. However, the estimated temperatures are impacted by the joint optimization of both modalities in the same MLP, as evidenced in Fig. 5 where the error in temperature prediction is high on the ROI. On the contrary, the average MAE and MAE for ThermoNerf mark a significant improvement over the other methods (for the MAE C on building scenes and C for others and for the MAE C for buildings and C for others).
@@
Additionally, we visually compare examples of test views rendered by each of the baseline methods as well as by ThermoNeRF (see Fig. 4). We observe noisy renderings with , while Nerfacto generates sensible reconstructions of the geometry. However, temperature predictions of Nerfacto are influenced by the RGB modality, leading to biased temperature values, sharper edges compared to the ground truth, and elements from RGB images that are visible in the synthesized thermal images. For example, in Fig. 4, for each Nerfactorgb+th image, the colors have leaked into the thermal images and buildings are warmer than the ground truth. While the building (spring) scene (top-row of Fig. 4) looks sharper when synthesized using Nerfactorgb+th compared to ThermoNeRF, in Fig. 5 top left row, one can see that the sharpness of the image comes with a high error rate on the temperature rendering—with a MAE of as opposed to for ThermoNeRF. Another example of detail leaking from the RGB into the thermal images can be seen in Fig. 5 and Fig. 6, where edges and structures are corresponding to high errors in Nerfactorgb+th.
@@
MAE=0.32°C
MAE=1.50°C
MAE=0.23°C
MAE=1.69°C
MAE=3.50°C
MAE=0.43°C
MAE=6.60°C
MAE=6.26°C
MAE=1.83°C
MAE=2.20°C
MAE=1.57°C
MAE=0.72°C
MAE=3.81°C
MAE=5.91°C
MAE=0.68°C
MAE=1.45°C
MAE=1.03°C
MAE=0.81°C
MAE=6.55°C
MAE=7.25°C
MAE=0.96°C
MAE=1.47°C
MAE=1.10°C
MAE=0.62°C
MAE=7.68°C
MAE=4.46°C
MAE=1.84°C
MAE=1.80°C
MAE=0.83°C
MAE=0.60°C
For in-detail qualitative evaluation of the results, the per-pixel absolute errors in temperature prediction is visualized in Fig. 5. One can see that the highest errors for all models are observed at the edges, due to the ghosting effect of thermal images leading to blurry edges. It should be noted that ThermoNeRF exhibits the best thermal prediction results while Nerfacto shows the worst thermal predictions. Nerfacto shows higher errors across the scenes than ThermoNeRF due to the contamination of the thermal information by the RGB modality.
Furthermore, cloudy skies are known to provide ideal conditions for infrared measurements outdoors, since the measured object is protected against solar radiation. This fact is supported by the comparison of the thermal image synthesis of Building Spring (top-left) and the Exhibition Building scenes. During measurements, the Building Spring scene’s sky is a mix of clear and cloudy images leading to more noise in the image synthesis than the second scene that consists of a uniform cloudy sky.
@@ 6.5 RGB View Synthesis
MED | Dorm 1 | Dorm 2 | INR | BI | Avg | ||||||
| PSNR | Nerfacto (RGB) | 19.88 | 20.02 | 22.18 | 17.53 | 19.13 | 19.11 | 14.17 | 20.38 | 19.06 | ||||||
| ThermoNeRF | 19.30 | 19.74 | 22.81 | 16.40 | 18.70 | 19.20 | 13.39 | 20.90 | 18.80 | |||||||
| Difference | -0.58 | -0.28 | +0.83 | -0.13 | -0.43 | +0.09 | -0.78 | +0.52 | -0.09 | |||||||
| SSIM | Nerfacto (RGB) | 0.64 | 0.62 | 0.64 | 0.43 | 0.57 | 0.61 | 0.39 | 0.65 | 0.57 | ||||||
| ThermoNeRF | 0.63 | 0.61 | 0.66 | 0.38 | 0.55 | 0.61 | 0.35 | 0.69 | 0.56 | |||||||
| Difference | -0.02 | -0.01 | +0.02 | -0.05 | -0.02 | +0.00 | -0.04 | +0.04 | -0.01 | |||||||
| LPIPS | Nerfacto (RGB) | 0.24 | 0.34 | 0.24 | 0.32 | 0.17 | 0.19 | 0.59 | 0.27 | 0.29 | ||||||
| ThermoNeRF | 0.27 | 0.34 | 0.24 | 0.38 | 0.19 | 0.22 | 0.61 | 0.28 | 0.32 | |||||||
| Difference | +0.03 | 0.00 | 0.00 | -0.06 | -0.02 | -0.03 | -0.02 | -0.01 | -0.01 |
| Metric | Method |
Trees | Laptop | Avg | ||||||||||||
| PSNR | Nerfacto (RGB) | 18.11 | 23.00 | 24.71 | 17.97 | 19.05 | 19.36 | 20.37 | 19.3 | 20.23 | ||||||||||||
| ThermoNeRF | 18.16 | 21.05 | 25.69 | 16.20 | 19.02 | 19.07 | 18.67 | 18,83 | 19.58 | |||||||||||||
| Difference | +0.05 | -1.95 | +0.98 | -1.77 | -0.03 | -0.29 | -1.61 | -0.47 | -0.63 | |||||||||||||
| SSIM | Nerfacto (RGB) | 0.53 | 0.75 | 0.73 | 0.52 | 0.65 | 0.73 | 0.67 | 0.61 | 0.65 | ||||||||||||
| ThermoNeRF | 0.56 | 0.60 | 0.77 | 0.43 | 0.65 | 0.72 | 0.59 | 0.61 | 0.62 | |||||||||||||
| Difference | +0.03 | -0.15 | +0.04 | -0.09 | 0.00 | -0.01 | -0.08 | 0 | -0.03 | |||||||||||||
| LPIPS | Nerfacto (RGB) | 0.25 | 0.23 | 0.42 | 0.17 | 0.28 | 0.18 | 0.32 | 0.21 | 0.26 | ||||||||||||
| ThermoNeRF | 0.25 | 0.24 | 0.37 | 0.18 | 0.30 | 0.19 | 0.35 | 0.22 | 0.26 | |||||||||||||
| Difference | 0.00 | +0.01 | -0.05 | +0.01 | +0.02 | +0.01 | +0.03 | +0.01 | 0.005 |
We investigate whether there is a significant degradation in the quality of RGB views compared with the standard Nerfacto trained only on the RGB inputs. Table 4 and Table 5 summarize the results, showing little to no degradation in the quality of the RGB views. Indeed, the differences between the PSNR, SSIM and LPIPS of Nerfacto and ThermoNeRF in RGB are respectively , and on building scenes, and , and on other scenes. The results support that using ThermoNeRF leads to higher accuracy in temperature without a significant degradation in the quality of the reconstructed RGB views. Thus, our approach effectively learns to model 3D scenes across both RGB and thermal modalities, making it suitable for various downstream tasks including building retrofit, energy consumption analysis and non-destructive infrastructure inspection.
@@ 7 Conclusion
In this work, we propose ThermoNeRF, a novel multimodal approach that leverages NeRF to render novel RGB and thermal views of a scene. Our method facilitates thermal analysis of buildings by using a sparse set of images captured from various angles around the structure. We curated a new dataset specifically tailored for RGB+thermal scene reconstruction, and our experimental results show that ThermoNeRF excels at synthesizing thermal images. It achieves an average MAE of 0.41°C on indoor scenes and 1.13°C on buildings, representing respective improvements of over 55% and 48% compared to the baseline approach that concatenates RGB and thermal modalities. A potential limitation of our method is its reliance on paired RGB and thermal images, as collecting such data requires precise camera calibration to ensure accurate alignment between the two modalities, which can pose practical challenges. To address this, future work will focus on developing techniques to train ThermoNeRF using unpaired RGB and thermal image data. We believe our method paves the way for new opportunities in thermal building reconstruction. For instance, future models could analyze ThermoNeRF’s outputs to detect failures in HVAC systems, piping, or electrical networks. We leave these exciting possibilities for future exploration.
Acknowledgements
We sincerely thank Jiarui Yu and Juliette Parchet for helping with the data collection and for reporting the results of ThermoNeRF on those datasets. We would also like to thank Ali Waseem for running some of the experiments.
This research was supported by Innosuisse - Swiss Innovation Agency, Innosuisse Grant no 105.237.1 IP-ICT titled Insulated: integrated solution for lean and abridged thermal evaluation with digital twins in collaboration with Schindler AG.